OpenStudy (anonymous):

2C8H18+25O2 ->16CO2+18H2O I need help to work out the enthalpy for the combustion of octane....^^balanced equation is above. I keep doing something wrong which causes me to get really high numbers. Please help.....

OpenStudy (chmvijay):

2C8H18+25O2 ->16CO2+18H2O Left 16=C 36=H 50=O Right 36H 32+18=50 =O 16=C it s all right :)

OpenStudy (anonymous):

i know that the balancing is right. I need help to work out the enthalpy using these bond energies. C-H: 413, C-C: 347, O=O: 498, C=O: 799, H-O: 464.

OpenStudy (chmvijay):

yes find the bond energy of the right molecules using above formula for example there X no of C-H bond in octane so X*413 = P and there are Y no of C-C bonds so = Y*347 = Q P+q will give you the energy of that molecule similarly you find for all and then energy of right molecules - energy of left molecule;es you do u get ur results

OpenStudy (chmvijay):

did u get what i am trying to say you ?

OpenStudy (chmvijay):

@adriana_98 ??????

OpenStudy (anonymous):

@adriana_98 are you there?

OpenStudy (anonymous):

Yes i am here @chmvijay @Jesstho.-.

OpenStudy (anonymous):

2C8H18+25O2 ->16CO2+18H2O 2[(18xC-H) + (7xC-C)] + 25(O=O) – 16(4xC=O) + 18(2xH-O) 2(7434 + 2429) +12450 – 16(3196) + 18(928) (19 544+ 12 450) – (51 136 + 16 704) 31 994 – 67 840 = -35 846 kJ The number is really big and when i divide it by two to get 1 mL of octane,i get -17 923 kJ. It just seems too big.

OpenStudy (chmvijay):

final answer is for 2 mole of octane combustion u dived it by 2 you get the combustion of octane per mole

OpenStudy (anonymous):

sorry, i am a bit confused??

OpenStudy (chmvijay):

2C8H18+25O2 ->16CO2+18H2O This is for 2 moles of C8H18 u have to find for 1 mole of C8H18 right?...

OpenStudy (anonymous):

yes, is it correct that i need to divide my final answer by 2??

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

:) then what is the problem ur facing then ?

OpenStudy (anonymous):

the answer is really big and i am not sure if it is right. Is the working out correct?

OpenStudy (anonymous):

And also another question, how do i find enthalpy for blends of fuels so i need to find it for 1 part octane and 9 parts ethanol. What process do i need to take to find that and is there a equation i can use aswell. I am not sure

OpenStudy (chmvijay):

its correct u jhave made one mistake in 16(4xC=O) + 18(2xH-O) not this but 16(2xC=O) - 18(2xH-O) i just did it and got 10,096 if i didved by 2 i get 5048 KJ/mol which can be compareble with real values as shown in this table http://www.ausetute.com.au/heatcomb.html

OpenStudy (chmvijay):

2[(18xC-H) + (7xC-C)] + 25(O=O) – 16(4xC=O) + 18(2xH-O) 12450+14868+4858-(25568+16704)=32176-42272 = -10096 KJ

OpenStudy (chmvijay):

look at once again ur calcualtions :P

OpenStudy (anonymous):

what bond energies did you use to do your calculations 2 messages earlier because i can't work out how you got those answers. Sorry for all the questions

OpenStudy (chmvijay):

the one you have mentioned above :)

OpenStudy (anonymous):

Nevermind i got the same answer as you

OpenStudy (chmvijay):

really ?....

OpenStudy (anonymous):

yes,i mucked around with it

OpenStudy (anonymous):

now i am unsure how to do this for blends. Do you know??

OpenStudy (chmvijay):

thats good :)

OpenStudy (chmvijay):

any other information they have given for that ?

OpenStudy (anonymous):

no, nothing

OpenStudy (anonymous):

the blends are like 1 part octane, 9 parts ethanol.... 2 parts octane, 8 parts ethanol etc... until we get to 0 parts octane and 10 parts ethanol. If that helps....

OpenStudy (chmvijay):

u mean to find enthalpy for blends of fuels ;)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i did a blend of 1 mL so for example for 1:9, i used 0.1mL of octane and 0.9mL of ethanol. I just need to know how to work out the enthalpy in mL instead of moles this time. If you know how to do that...

OpenStudy (chmvijay):

 find our enthalapy of the combustion for one mole of ethanol first similarly as we did above:)